OpenStudy (anonymous):

9y - 11 + 7 ?? how do i do this?

OpenStudy (anonymous):

Put your varibles on one side of the = sign and your constants on the other. then you must divide both sides by 9 inorder to give you a statment y=?

OpenStudy (amistre64):

there isnt an equals sign so that would be rather pointless :)

OpenStudy (amistre64):

add the like terms and leave it as is....

OpenStudy (anonymous):

9y=-11+7 9y=16

OpenStudy (anonymous):

am i wrong? lol

OpenStudy (amistre64):

9y does not equal -11+7

OpenStudy (amistre64):

if anything: 9y -11+7 = 9y -4

OpenStudy (anonymous):

is there a equal sign in your problem?

OpenStudy (amistre64):

yes, and heres why. 5 = 5 is a true statement 8r=8r is a true statement 9y-4 = 9y-4 is a true statement solving for y? it can be any number it wants to be...thers nothing restricting its use right?

OpenStudy (amistre64):

9y-4 = 0 can be solved for y IF and only IF 9y-4 = 0

OpenStudy (amistre64):

otherwise: 9y-4 = 9y-4 its an identity

OpenStudy (amistre64):

barring any typos from the poster :) theres not much else we can do with it

OpenStudy (anonymous):

I agree I was just putting in a = sign becaue i thought she was trying to solve for y.

OpenStudy (amistre64):

I understand :) and that would have been a good approach, except for we just dont know what to equal it to to begin with ;)

OpenStudy (anonymous):

amistre64 can you a take a look at the question i posted about the tangent line please

OpenStudy (amistre64):

i can if I can find it :) is it recent or past?

OpenStudy (anonymous):

about 15 min ago, i will repost